Calculate the Producers Surplus if the supply equation is p = 7+ 2q1/3 and the unit price is set at p = 13.
Answer #1

P = 7+292. © = 13 13=77@gs 29*=13-7. 293-6. q=33: 9-27 Producer sapluus = pq- da Pa $(7)= 7+29s. RS = (13/07) - JChaq) de